OpenStudy (chemicaldiscoboy):

Which point is on the graph of y=3/x? 6,2 6, 1/2 0,3 3,0

OpenStudy (ajspeller):

Substitute in the values of x and y into the equation y=3/x. Whichever one is true is the answer.

OpenStudy (pnathaniel56):

y=3\x hsve u guesed yet

OpenStudy (ajspeller):

example for (6,2), 2=3/6, 2=1/2 not the correct answer. Keep testing the points.

OpenStudy (pnathaniel56):

i would go with3,0

OpenStudy (chemicaldiscoboy):

I did 6,2 because 3 x 2 = 6. Is that how you do it?

OpenStudy (chemicaldiscoboy):

@ajspeller

OpenStudy (chemicaldiscoboy):

The answer was 6, 1/2

OpenStudy (mathmale):

Which point is on the graph of y=3/x? (6,2) (please enclose these coordinates in parentheses) Here x=6. Substitute 6 for x in y=3/x. Is your result equal to 2, or is it not? If it is not, then (6,2) is not a solution of the given equation. Please do the other 3 in the same way.